OpenStudy (anonymous):

You roll two dice. What is the probability that the sum of the dice is even or the sum of the dice is greater than 6? Make sure you only account for dice sums that are both even and greater than 6 once. A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

Total number of possible outcomes = 6^2 = 36 Sum of the digits---No. of outcomes 02 ------------------------- 01 03 ------------------------- 02 04 ------------------------- 03 05 ------------------------- 04 06 ------------------------- 05 07 ------------------------- 06 08 ------------------------- 05 09 ------------------------- 04 10 ------------------------- 03 11 ------------------------- 02 12 ------------------------- 01 Total --------------------- 36 P(sum of the digits is odd) = P(A) = (2+4+6+4+2) / 36 = 18/36 P(sum of the digits > 6) = P(B) = (6+5+4+3+2+1) / 36 = 21/36 P(A and B) = (6+4+2) / 36 = 12/36 Required probability = P(A or B) = P(A)+P(B)-P(A and B) =18/36 + 21/36 - 12/36 = 27/36 = 3/4

OpenStudy (anonymous):

you're welcome.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

did it help?